how many atoms of copper are there in a piece of copper that weighs the same as a piece of aluminum that contains 4.86 x 10(21) atoms of aluminum
1
Expert's answer
2019-05-03T07:19:49-0400

Let's find out the amount of Aluminium.

n = NAl / NA = 4.86 × 1021 / 6,02 × 1023 = 0.807 × 10-2 (mol);

Now we can find out the mass of Aluminium:

mAl = n × Mr = 0.807 × 10-2 × 27 = 0.218 (g);

As the masses of Aluminium and Copper are equal, we can calculate the amount of Copper:

n = mCu / Mr = 0.218/64 = 0.0034 (mol);

NCu = n × NA = 0.0034 × 6.02 × 1023 = 2.05 × 1021